When the length of each side of a square is increased by 5 inches, the area of the resulting square is 2.25 times the area of the original square. What is the area of the original square?
Answer by edjones(8007)

Let s=original side
A=s^2
.
(s+5)^2=2.25s^2
s^2+10s+25=2.25s^2
1.25s^2-10s-25=0 subtract the left side from the right.
s^2-8s-20=0 divide each side by 1.25
(s+2)(s-10)=0
s=10
10^2=100 sq in. area of the original square.
.
Ed